### v1

# ShaRP: SHAllow-LayeR Pruning for Video Large Language Models Acceleration

Abstract
Video Large Language Models (VLLMs) face the challenge of high computational load during the pre-filling stage due to the processing of an enormous number of visual tokens. Although attention-based pruning methods are widely used to accelerate inference, trials at early decoder layers often result in significant performance degradation, especially under high compression rates. We argue that while attention-based pruning inherently holds the potential to identify the most relevant visual tokens, its effectiveness in shallow decoder layers is limited by factors such as positional encoding bias and insufficient information interaction. In this paper, we propose an improved attention-based pruning framework, termed ShaRP, that integrates segment-aware causal masking, positional debiasing, and token deduplication for enhanced token selection. It enables effective pruning at shallow layers while maintaining stable performance under high compression rates without retraining. Extensive experiments demonstrate that ShaRP achieves competitive performance across multiple video understanding benchmarks, establishing a new paradigm for accelerating VLLM inference.
1 Introduction
Vision Large Language Models (VLLMs) [1, 2, 3, 4, 5] have brought deep video understanding task into a new era. They use a vision encoder to convert video streams into many tokens, which interact during pre-filling. Such a process thus incurs quadratic computational complexity relative to the token count. To mitigate this, numerous studies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] have focused on reducing visual token counts to compress information and accelerate VLLM inference.
Among various compression strategies, attention-based ones [6, 7, 8, 21, 14] offer a compelling approach. Unlike visual cue-guided pruning that operates purely on visual similarity [10, 22, 23, 24, 13], attention-based methods leverage the inherent cross-modal reasoning ability of the model to estimate token importance, inherently aligning the pruning process with semantic relevance. It not only enhances interpretability [25, 26, 27] but also enables adaptive compression that dynamically focuses on content most relevant to the textual query. Representative studies such as FastV [6] show that visual information becomes less critical in deeper decoder layers, motivating aggressive pruning early. However, this strategy struggles with high compression on long videos, as performance drops sharply, leading subsequent works [7, 8, 19] to prune more conservatively in deeper layers, losing much of the computational savings. Viewed differently, we advocate the potential of attention in shallow layers as an effective criterion for token selection. To explore this, we conduct a pilot study comparing attention scores across shallow and deep layers. As shown in Fig. 2, visual analysis reveals that visual tokens near the sequence end receive abnormally high attention in shallow layers. When properly tackled, shallow layer attention exhibits distributions similar to deep ones, enabling early yet reliable pruning.
In our paper, we contend that early-stage pruning failure is not inevitable; it can be achieved at an early stage (Fig. 1) by addressing three overlooked factors:
(1) Information remains under-interacted in shallow layers: As revealed in HiMAP [14] and visiPurner [28], shallow layers mainly facilitate cross-modal exchange, while interactions among same-modality visual tokens remain underdeveloped. Once pruning at shallow layers, some retained visual tokens often exhibit high-norm anomalies [29]; they receive high attention scores yet carry low information density (i.e., attention collapse in [30, 31]). This frequently occurs with semantically irrelevant tokens, such as background visual tokens. This issue arises because the softmax function assigns non-zero attention to all tokens, causing many low-relevance or non-priority tokens to receive unnecessary attention.
(2) Content-agnostic positional biased can be introduced [32]: Such insufficient interaction, as discussed in (1), leaves shallow-layer attention overly dependent on positional priors. Taking RoPE [33] as an example, due to its long-term decay property, the model assigns unreasonably high scores to tokens near the end of the sequence (i.e., to later video frames). In methods like FastV [6], when the compression ratio increases (e.g., pruning over 50% tokens), most retained tokens come from the end of the sequence. This inherent bias in attention scores makes attention-based pruning in shallow layers unreliable.
(3) Post-Pruning stage faces an importance-redundancy dilemma: Even after mitigating positional bias through methods like (2), relying solely on attention as an importance indicator can still cause the model to repeatedly retain high-attention [29] tokens from adjacent frames or similar backgrounds. As observed in the first layer of LLaVA-Video-7B [2], over half of the tokens selected based on attention scores exhibit high redundancy [24], with cosine similarity exceeding 0.9 to other selected tokens. Under a constrained compression rate, these similar register tokens [29] "crowd out" those that should preserve the holistic context of the scene, ultimately leading to degraded performance under high compression ratios.
Our work is built upon a core insight: effective attention-based pruning in shallow layers is achievable when the three limitations above are properly addressed. To this end, we propose a novel training-free framework comprising three components — Segment-Aware Causal Masking (SegM), Positional Bias Calibration (PosC), and Register Token Deduplication (RegD) — for efficient shallow-layer token pruning in video understanding. First, to enhance information interaction, inspired by [34], which observed that locally focused heads in Transformers form near-diagonal block attention, segmenting complex reasoning into phrasal units. We adapt this to the visual domain by partitioning videos into content-aware segments, intensifying local interactions and alleviating attention collapse [30, 31]. Second, our method calibrates positional bias in attention scores through a rapid global bias estimation. Finally, to balance between the importance and redundancy, high-attention register tokens are refined through a three-step pipeline composed of pre-filtering, deduplication, and post-filling, which eliminates redundancies while enriching the representation with broader contextual information. Extensive experiments demonstrate that our approach achieves high compression in shallow LLM layers while maintaining state-of-the-art performance, establishing a new paradigm for attention-based token pruning.
Our contributions are summarized as follows:
Revealing the Factors of Early Pruning Failure: We show that attention allocation collapse, PE bias, and register token redundancy are three factors that limit attention-based pruning performance in shallow decoder layers.
A Simple yet Effective Training-Free Framework: We introduce a plug-and-play framework, ShaRP, integrating segment-aware causal masking, positional bias calibration and register token deduplication to effectively address the above factors and improve compression performance.
2 Related work
2.1 Visual Cue-Guided Token Pruning
Since their emergence, VLLMs [35, 36, 37, 3, 38, 39, 4] have demonstrated significant potential across diverse applications, including VQA [40, 41, 42], video captioning [43, 44], contextual reasoning [45, 46, 47] etc.. Earlier approaches employ a limited number of learnable queries to aggregate visual information. For instance, VideoLLaMA [48] adopts a Q-Former [39] to distill dense visual features into compact query sets. While GlimpsePrune [49] introduces a trainable Visual Importance Predictor (VIP) within the VLLM to score and prune visual tokens at different layers. The recent method EPIC [50] adopts a Progressive Consistency Distillation strategy to re-finetune the model, enabling it to learn rich visual information even from incomplete visual inputs. Token pruning can be achieved in a training-free manner. Among these, LLaVA-PruMerge [9], VisionZip [10], and VisPruner [12] leverage the inherent sensitivity of vision encoders to visual tokens to identify and retain visually dominant tokens. DivPrune [22] formulates pruning as a Max-Min Diversity Problem (MMDP), ensuring that the set of retained visual tokens possesses maximum diversity. Nevertheless, the temporal redundancy inherent in video data calls for temporal-aware pruning mechanisms. DyCoke [13] merges tokens within fixed temporal windows. Other methods [23, 18, 19] remap the numerous visual tokens back to the frame level, enabling them to identify and merge highly similar visual content across frames while retaining unique content. Among these, LLaVA-Scissor [23] identifies non-overlapping visual regions via Semantic Connected Components (SCCs). FastVID [18] performs segmentation based on video content, combining temporal segmentation with spatio-temporal merging. HoliTom [19] utilizes dynamic programming to merge redundant information across adjacent frames.
2.2 Text–Vision Attention-Based Token Pruning
As the other family of token pruning approaches, pruning visual tokens based on text–vision attention scores in the transformer layers of the LLM decoder has proven effective for reducing redundancy while preserving task-relevant information. For example, FastV [6] pioneers this path as it observes strong sparsity in visual-token attention at deeper layers and thus prunes a fixed proportion of tokens based on shallow-layer attention scores. However, this approach is susceptible to positional bias and attention collapse, which inhibits high compression ratios in shallow layers. To achieve smoother compression dynamics, PyramidDrop (PDrop) [7] adopts a hierarchical pruning mechanism while Feather [8] introduces a de-RoPE-based attention criterion and performs multi-stage pruning via uniform sampling, yet the bias from the positional encoding at earlier layers and from the visual encoder remains in the de-RoPEed attention scores. HoloV [31] mitigates attention collapse by pruning tokens within uniform spatial partitions based on contextual diversity and saliency. However, this block-wise strategy may scatter attention across irrelevant regions, weakening focus on question-relevant content in long videos. HiMAP [14] adopts a layer-adaptive strategy that prunes tokens with low cross-modal contribution in shallow layers and those less relevant to intra-visual aggregation in deeper layers. In addition, FrameFusion [24] performs layer-wise compression in the shallow layers of LMMs by exploiting the similarity among visual tokens, while applying attention-based importance pruning to the remaining tokens in deeper layers. However, relying solely on attention scores can overemphasize a few salient tokens, losing holistic context and risking semantic distortion when merging dissimilar tokens. This underscores the need for both reliable attention evaluation and careful token merging to preserve meaningful information.
3 Preliminary Knowledge
Multimodal Causal Attention: VLLMs such as LLaVA [3] and Qwen [4] process both encoded visual and text tokens with causal attention mechanism. Given a visual token set and a textual token set , their tokens are concatenated into with the length of sequence and the hidden dimension, and iteratively processed within decoder layer by layer, via the causal attention mechanism.
Specifically, is first linearly projected into query , key , and value matrices, each in . The attention score matrix and output are computed as:
where is a causal mask with entries for all and are token indices, to effectively mask the upper triangular region. This ensures that each token only attends to the past tokens.
Rotary Positional Encoding: RoPE [33] encodes relative position through rotary transformations. Instead of using additive positional biases, it applies position-specific rotary matrices to the query and key at positions and . The resulting relative position-aware attention score based on Eq. 1 is thus given by:
where denotes relative distance between and .
4 Methodology
Current attention-based pruning methods struggle in early decoder layers due to attention allocation collapse, positional encoding bias, and register token redundancy. As illustrated in Fig. 3, we propose a training-free framework that respectively addresses these issues through effective yet simple techniques, detailed as follows:
4.1 Segment-Aware Causal Masking
In the shallow layers of large video-language models, high-norm but semantically weak tokens often emerge, typically corresponding to redundant or low-information regions [29, 10]. Relying solely on attention scores for importance-based pruning exacerbates this issue, as the causal attention and softmax normalization mechanisms inherently assign non-zero weights to all tokens, allowing even irrelevant ones to receive disproportionately high attention [29]. Meanwhile, the insufficient interaction among visual tokens at early layers [14] further limits the model’s ability to aggregate meaningful representations, leaving these anomalous tokens unrefined.
To mitigate this attention collapse [30, 31] that emerges in shallow layers, we draw inspiration from locally focused heads [34], which concentrate attention within small, semantically coherent regions. We thus introduce segment-aware causal masking that constrains visual attention within content-consistent video segments. This localized masking preserves the causal decoding property while enhancing intra-segment visual interaction, ensuring that information flows more effectively within relevant visual contexts.
Precisely, given the visual token set , we first apply average pooling to tokens from the same frame, obtaining a frame-wise averaged token set , where is the frame index. To partition into semantically consistent segments, we insert segment boundaries by evaluating the cosine similarity [51] between consecutive tokens in against a predefined threshold :
where is the size of . The boundaries help us to split into contiguous non-overlapping segments . Based on , we then perform a masked attention computation where visual tokens within the same segment remain mutually visible, while those from different segments are completely isolated. Such segment-aware visual mask is thus constructed as a block-diagonal matrix as show in Fig. 3(a). The segment-aware masked attention scores between one visual token to others follows:
which implies that attentions originally scattered across all segments will be squeezed into one single segment. This localized masking effectively narrows the operational range of the softmax, concentrating attention weights within semantically consistent regions (i.e, smaller triangles) and enhancing the information flow of visual feature aggregation.
4.2 Positional Bias Calibration
Positional Encodings (PE) equip VLLMs with token order information, whereas also introduce inherent positional bias [53] that favors tokens at specific locations. A typical example is the long-term decay issue [30] in RoPE (Eq. 2), where attention scores decrease as the relative token distance increases. While in the context of video token pruning, existing methods [6, 7, 8] usually identify more important visual tokens by referring attention scores from the last text token . They typically assume that visual tokens with higher scores are more relevant to the query. However, we argue that positional bias significantly distorts the distribution of attention scores, which may lead to misinterpretation of this mechanism. As shown in Fig. 4 (left-top), our analysis on the VideoMME [52] benchmark at 20% retention reveals that most top-ranked tokens (orange regions) cluster near the final frames, reflecting a clear positional bias rather than true semantic relevance. Fig. 4 (right-middle) further illustrates a case where the question targets a mid-video event, yet attention remains skewed toward later frames, leaving semantically important regions under-attended (blue regions).
Therefore, we propose to calibrate PE bias before pruning them with respect to their attention scores. Notably, PE techniques span a broad spectrum, ranging from Absolute PEs (APE) [54, 55, 56] to Relative ones (RPE) [33, 57, 58], with each exhibiting distinct properties (e.g., additivity, rotation invariance) and bias patterns. Furthermore, as VLLMs may employ multiple PE strategies for different modules (e.g., visual encoder, text encoder, decoder), simply removing one [8] or all PEs for debiasing will cause the remaining PE components to dominate or even inference collapse.
Without loss of generality or effectiveness, and inspired by the additive formulation of relative position encoding [59], we propose a simple, content-agnostic, and PE-type-independent method to estimate the positional encoding bias inherent in VLLMs. To this end, we input a video of entirely black frames into the VLLM, ensuring all visual tokens are homogeneous and deprived of information. In this setup, any deviation in the attention score between and directly reflect the bias introduced by positional encodings. Larger values indicate stronger bias, requiring greater compensation on , and vice versa. To remove this bias, we simply subtract it from the attention scores:
where is a hyperparameter that controls the strength of the compensation. Note that this compensation on is only for token ranking and pruning, the value of each token remains unaltered. After applying our proposed debiasing strategy, as shown in Fig. 4 (left-bottom), the overall attention distribution becomes more balanced across the temporal dimension. Meanwhile, in the right-bottom panel, attention is more concentrated on semantically relevant regions (e.g., hats), effectively alleviating the positional bias issue. Importantly, this bias estimation requires only a single lightweight inference with bias-free inputs and is performed once per pruning step, while all subsequent attention computations remain fully compatible with FlashAttention [60].
4.3 Pruning and Register Token Deduplication
Pruning. After applying segment-aware causal masking (Sec. 4.1) and attention debiasing (Sec. 4.2), following prior works [6] we perform token pruning based on the debiased attention scores . Specifically, we select the top- visual tokens with the highest scores across the entire sequence. These retained tokens are distributed among different segments , and form a compact register token set .
Register Token Deduplication. Although high-attention tokens are preserved after pruning, many still remain highly similar in both spatial and temporal dimensions, as discussed in Sec. 1. To further compress the representation, we introduce a three-step post-pruning refinement for register tokens within each segment, consisting of pre-filtering, deduplication, and post-filling, taking register tokens within segment as an example, i.e. :
(1) Token Pre-Filtering: Non-register tokens whose similarity to any token in exceeds are absorbed into their nearest similar register token. The remaining non-register tokens (noted as ) are preserved for Step (3).
(2) Token Deduplication: We perform iterative merging on register tokens by sequentially comparing adjacent tokens. A token is absorbed into its preceding token if their similarity exceeds , otherwise it becomes the new pivot. This continues until the end of the sequence is reached, giving a deduplicated register token set for Step (3).
(3) Token Post-Filling: Step (2) reduces redundancy among register tokens by merging similar ones. To maintain a fixed number of register tokens under the preset compression rate, we cluster the non-register tokens in and replenish with selected cluster centers. Starting from the first token in , we perform iterative clustering: a subsequent token joins the current cluster if its similarity exceeds , otherwise it initiates a new cluster. For the resulting clusters, we compute the diversity score of each of their centers as:
where and denote set sizes and is a balancing hyperparameter. This score considers both the average similarity to register tokens and the cluster size. Centers with the highest diversity are selected as replenished register tokens. Finally, the filled register token set is forwarded to deeper decoder layers for inference.
| Method |
MVBench ↑ |
MLVU ↑ | VideoMME ↑ | Average ↑ | |||||||||
| Overall | Short | Medium | Long | |||||||||||||
| Duration | 16s | 160min | 3120min | 160min | 13min | 330min | 3060min | Score | % | |||||||
| LLaVA-OneVision [3] | 100% | 40.8 | 57.9 | 56.4 | 63.2 | 58.6 | 70.3 | 56.6 | 48.8 | 59.0 | 100 | |||||
| FastV [6] | 25% | 11 | 55.3 | 55.1 | 60.3 | 56.5 | 68.2 | 55.0 | 46.3 | 56.8 | 96.2 | |||||
| PDrop [7] | 28.5% | 11 | 56.7 | 54.2 | 61.1 | 56.4 | 68.0 | 53.6 | 47.7 | 57.1 | 96.7 | |||||
| Feather [8] | 25% | 11 | 54.8 | 54.7 | 60.8 | 55.0 | 65.3 | 53.6 | 46.2 | 56.3 | 95.4 | |||||
| FrameFusion [24] | 31% | - | 56.7 | 55.8 | 62.2 | 57.4 | 68.1 | 55.4 | 48.6 | 57.8 | 97.9 | |||||
| ShaRP | 28.2% | 11 | 57.0 | 56.2 | 62.7 | 57.9 | 69.8 | 56.9 | 48.0 | 58.5 | 99.0 | |||||
| FastV [6] | 20% | 9.31 | 55.8 | 55.3 | 59.5 | 55.8 | 66.4 | 54.2 | 46.8 | 56.6 | 95.9 | |||||
| PDrop [7] | 19.6% | 9.31 | 55.4 | 51.8 | 59.9 | 55.7 | 66.9 | 53.3 | 46.9 | 55.7 | 94.4 | |||||
| Feather [8] | 20% | 9.31 | 54.3 | 54.8 | 59.8 | 54.8 | 65.1 | 53.6 | 45.6 | 55.9 | 94.7 | |||||
| FrameFusion [24] | 26% | - | 56.0 | 55.1 | 61.7 | 57.5 | 68.0 | 55.9 | 48.6 | 57.6 | 97.5 | |||||
| ShaRP | 23.5% | 9.31 | 57.1 | 56.1 | 62.4 | 58.0 | 68.5 | 56.1 | 48.6 | 58.4 | 98.9 | |||||
| FastV [6] | 15% | 7.66 | 54.9 | 54.8 | 58.4 | 54.7 | 65.1 | 53.4 | 45.7 | 55.7 | 94.4 | |||||
| PDrop [7] | 11.5% | 7.66 | 54.5 | 50.0 | 59.1 | 54.0 | 64.3 | 51.6 | 46.1 | 54.4 | 92.2 | |||||
| Feather [8] | 15% | 7.66 | 54.5 | 53.0 | 59.1 | 54.7 | 65.1 | 53.3 | 45.6 | 55.3 | 93.7 | |||||
| FrameFusion [24] | 21% | - | 57.0 | 54.3 | 60.6 | 56.6 | 67.0 | 55.2 | 47.6 | 57.1 | 96.8 | |||||
| ShaRP | 18.8% | 7.66 | 56.3 | 55.7 | 61.6 | 57.4 | 68.1 | 56.7 | 47.5 | 57.8 | 97.9 | |||||
| FastV [6] | 10% | 6.03 | 53.5 | 52.4 | 56.2 | 52.6 | 61.1 | 51.6 | 45.2 | 53.7 | 90.9 | |||||
| PDrop [7] | 1.5% | 6.03 | 48.2 | 44.7 | 52.5 | 45.0 | 49.6 | 45.8 | 39.6 | 47.6 | 80.6 | |||||
| Feather [8] | 10% | 6.03 | 53.6 | 52.7 | 57.6 | 53.6 | 62.4 | 52.7 | 45.6 | 54.4 | 92.1 | |||||
| FrameFusion [24] | 16% | - | 55.1 | 53.0 | 58.3 | 55.5 | 65.8 | 54.1 | 46.7 | 55.0 | 94.0 | |||||
| ShaRP | 14.0% | 6.03 | 55.9 | 55.8 | 61.4 | 56.4 | 65.9 | 56.3 | 46.9 | 57.4 | 97.2 | |||||
| Method |
MVBench ↑ |
VideoMME ↑ | Average ↑ | ||||||||||
| Overall | Short | Medium | Long | ||||||||||||||
| Duration | 16s | 160min | 160min | 13min | 330min | 3060min | Score | % | |||||||||
| Qwen2-VL [4] | 2880 | 100% | 16.7 | 64.8 | 55.1 | 56.7 | 67.4 | 54.2 | 48.3 | 58.9 | 100 | ||||||
| FastV [6] | 2880/576 | 20% | 4.1 | 57.7 | 48.0 | 52.6 | 61.9 | 49.2 | 46.7 | 52.8 | 89.6 | ||||||
| PDrop [7] | 2880/662 | 23.1% | 4.1 | 54.4 | 48.2 | 52.0 | 61.4 | 48.9 | 45.8 | 51.5 | 87.6 | ||||||
| Feather [8] | 2880/576 | 20% | 4.1 | 60.2 | 50.0 | 53.0 | 62.1 | 50.6 | 46.4 | 54.4 | 92.4 | ||||||
| FrameFusion [24] | 2880/740 | 25.7% | - | 60.4 | 48.5 | 53.4 | 62.4 | 50.7 | 47.2 | 54.1 | 91.9 | ||||||
| ShaRP | 2880/679 | 23.6% | 4.1 | 64.5 | 52.4 | 53.0 | 62.4 | 50.2 | 46.5 | 56.6 | 96.2 | ||||||
| FastV [6] | 2880/288 | 10% | 2.6 | 52.2 | 45.6 | 50.9 | 57.4 | 48.4 | 46.7 | 49.6 | 84.2 | ||||||
| PDrop [7] | 2880/78 | 3.4% | 2.6 | 46.1 | 45.3 | 46.6 | 52.8 | 44.3 | 42.6 | 46.0 | 78.1 | ||||||
| Feather [8] | 2880/288 | 10% | 2.6 | 56.8 | 48.6 | 51.1 | 60.2 | 47.9 | 45.1 | 52.2 | 88.6 | ||||||
| FrameFusion [24] | 2880/472 | 16.4% | - | 59.5 | 47.4 | 52.0 | 61.3 | 49.0 | 45.8 | 53.0 | 90.0 | ||||||
| ShaRP | 2880/400 | 13.9% | 2.6 | 63.9 | 53.3 | 52.2 | 61.5 | 49.1 | 46.1 | 56.5 | 95.9 | ||||||
5 Experiments
5.1 Experimental Settings
Benchmarks. We evaluate ShaRP on four widely used video understanding benchmarks: MVBench [61, 62], LongVideoBench [63], MLVU [64], and VideoMME [52]. These datasets contain videos from 10 seconds to 2 hours, with over 10,000 questions spanning diverse scenarios, providing a solid test of our method’s robustness and generality.
Baselines. We compare our method with six representative training-free token pruning approaches. FastV [6] and PDrop [7] perform attention-guided pruning starting from the 2nd and 4th decoder layers, respectively. Feather [8] removes positional embeddings at the 2nd layer and applies uniform token sampling before pruning. FrameFusion [24] conducts progressive, layer-wise pruning based on token similarity and selects important tokens according to attention scores. To verify the compatibility of our strategy, we further integrate it into two representative pruning methods, VisionZip [10] and HoliTom [19]. Specifically, VisionZip performs outer-LMM frame-level pruning by identifying important tokens with a vision encoder, on top of which we apply our inner-LMM compression for further token reduction. HoliTom [19] adopts a two-stage pruning framework, where redundant tokens are progressively removed through outer- and inner-LMM pruning. We replace its inner-LMM pruning stage with our proposed strategy. For layer-wise compression methods, FrameFusion [24], we compute an equivalent retention ratio based on the average proportion of retained tokens across all LLM layers to match overall computational cost. Unless otherwise specified, we follow the official configurations of both backbones.
Implementation Details. We implement ShaRP on two representative multimodal LLMs, i.e. LLaVA-OneVision-7B [3] and Qwen2-VL-7B [4], which embody distinct architectural designs to ensure the robustness of our approach and we apply pruning at the first decoder layer. All experiments are conducted on NVIDIA A40 GPUs using the LMMs-Eval framework [65]. The hyperparameters are empirically determined and fixed to , , for all experiments. See more details in the supplementary materials.
5.2 Main Results
Enhanced Overall Performance. For LLaVA-OneVision [3], tokens from 32 sampled frames are evaluated under retention ratios about , with FLOPs matched across methods for fair comparison. Tab. 1 reports average scores and accuracy under different and TFLOPs. FastV [6] achieves decent performance at higher retention ratios (e.g., ), yet its accuracy declines sharply under aggressive pruning (e.g., ), indicating that raw attention scores without correction are unreliable when few tokens remain. Feather [8] alleviates positional bias by removing local positional encoding (at a certain layer) and uniform sampling. It shows limited performance under low retention ratios as many tokens retained are irrelevant to the query, achieving only 95.4% of the vanilla model’s performance even at 25% retention. PDrop [7] employs a progressive, layered pruning method. To achieve comparable FLOPs, this strategy necessitates aggressive pruning in deeper layers, which leads to very low accuracy even if the shallow layers retain the complete set of tokens. In contrast, ShaRP preserves up to 99% of vanilla accuracy even with aggressive shallow-layer pruning (72%). Its advantage is particularly pronounced on long-form video benchmarks such as LongVideoBench [63] and MLVU [64], where it surpasses the second-best method by 2.8% and 3.1% at 10% retention, respectively. These gains arise from ShaRP’s ability to aggregate information across temporally distant frames. Across all retention ratios, ShaRP consistently achieves competitive performance, demonstrating its robustness and adaptability.
Cross-Backbone Generalization. To evaluate the cross-model generalization of our approach, we extend experiments to Qwen2-VL [4], as detailed in Tab. 2. For this model, 32 sampled frames generate a token sequence as input, and we test with retention ratios across MVBench [61, 62], LongVideoBench [63], and VideoMME [52]. Unlike the LLaVA series, Qwen2-VL processes video frames using 3D convolutions and uses an MLP to aggregate tokens. For fair comparison under reasonable memory constraints, all methods are evaluated at the same input resolution. As shown in Tab 2, ShaRP consistently outperforms all baselines across the evaluated retention ratios, demonstrating its robustness across different backbones.
Compatibility Evaluation. We further test ShaRP’s compatibility by integrating it into two representative pruning frameworks, VisionZip [10] and HoliTom [19], covering both outer-LLM and two-stage paradigms. As summarized in Tab. 3, ShaRP consistently enhances both frameworks’ performance under equal FLOPs, confirming its compatibility.
Inference Efficiency. Tab. 4 presents the inference efficiency of ShaRP on the LLaVA-OneVision [3] benchmark. Our approach maintains strong performance while significantly reducing computational complexity. For instance, at a retention rate of 22.8%, TFLOPS decreases to 9.31, cutting pre-filling latency from 1181.6 ms to 538.5 ms (about 2.2 speedup). With a more aggressive rate of 14.7%, latency is further reduced to 393.4 ms, achieving approximately 3.0 acceleration. We further analyze the pre-filling time distribution across stages: debiasing + local masking + pruning (denoted as "Attn." in Tab. 4), token pre-filtering ("Fltr."), deduplication ("Dedpl."), post-filling ("Fil."), and the standard VLLM forward pass ("Fw."). Results show that debiasing and masking incur minimal overhead, whereas token deduplication accounts for nearly half of the total time, which highlights a key target for future optimization.
Visualization Comparison. We further perform visualization analysis in Fig. 5 (left). As shown, FastV [6] suffers from positional bias, omitting important visual segments, whereas Feather [8]’s partial debiasing and uniform sampling are less robust to scene transitions in long video QA tasks. In contrast, ShaRP ensures comprehensive coverage and adapts to temporal content variations, producing semantically coherent token selection even under aggressive compression. Additional examples can be found in the supplementary material.
MVBench ↑ |
MLVU ↑ | VideoMME ↑ |
| VisionZip [10] | 12.8% | 4.3 | 55.6 | 51.6 | 60.9 | 55.4 | 55.9 | ||||||||
[…]
| HoliTom [19] | 15%/7.5% | 4.3 | 58.4 | 55.7 | 62.1 | 57.2 | 58.4 | ||||||||
[…]
| TFLOPs | Prefilling Time (ms) ↓ | Average ↑ | |||||
| Attn. ↓ | Fltr. ↓ | Dedpl. ↓ | Fil. ↓ | LLM Fw. ↓ | Total ↓ | ||
| 40.8 (100%) | – | – | – | – | 1181.6 | 1181.6 (1.0) | 59.0 (100%) |
| 9.31 (22.8%) | 26.3 | 13.1 | 234.7 | 54.0 | 212.6 | 538.5 (2.2) | 58.4 (98.9%) |
| 6.03 (14.7%) | 26.3 | 12.1 | 143.6 | 52.0 | 159.4 | 393.4 (3.0) | 57.4 (97.2%) |
5.3 Ablation Study
Ablation and Synergy Analysis. We conduct ablation experiments on our three modules—Segment-Aware Causal Masking (SegM, Sec. 4.1), Positional Bias Calibration (PosC, Sec. 4.2), and Register Token Deduplication (RegD, Sec. 4.3)—under a 23.5% token retention ratio, with results in Tab. 5. PosC alone improves the average score from 55.8 to 57.4 by enhancing focus on critical visual regions. Adding SegM (PosC+SegM) further increases performance to 57.6, mitigating attention collapse and stabilizing temporal-contextual aggregation. Combining PosC with RegD raises the score to 58.0, demonstrating RegD’s effectiveness in compressing redundant tokens and enriching structural representation. Even without PosC, SegM+RegD reaches 57.6, confirming their independent contributions. Integrating all three modules achieves the best score of 58.4, validating the complementary benefits of each component
| Method | MVBench ↑ |
MLVU ↑ | VideoMME ↑ |
| Original | 54.4 | 53.0 | 61.0 | 54.7 | 55.8 | ||||
| PosC | 56.6 | 55.3 | 60.8 | 57.0 | 57.4 | ||||
| PosC+SegM | 56.6 | 55.5 | 60.9 | 57.2 | 57.6 | ||||
| PosC+RegD | 56.6 | 56.0 | 62.5 | 57.0 | 58.0 | ||||
| SegM+RegD | 56.5 | 55.0 | 62.2 | 56.7 | 57.6 | ||||
| PosC+SegM+RegD (ShaRP) | 57.1 | 56.1 | 62.4 | 58.0 | 58.4 |
Visualization and Analysis. We visualize the token selection results of the ablation study in Fig. 5 (right) to further analyze the effect of each component. The visualization shows that PosC enables the model to capture a broader and more semantically aligned visual context, while SegM and RegD encourage more diverse and informative token selection by reducing local redundancy.
6 Conclusion
In this work, we propose ShaRP, a training-free framework addressing the critical challenge of shallow-layer pruning in Video Large Language Models (VLLMs). We identify three key factors causing attention-based pruning to fail in early layers: attention collapse, positional encoding bias, and register token redundancy. To tackle these limitations in shallow-layer attention, we introduce a synergistic three-stage pipeline that enhances feature aggregation, calibrates attention distributions and removes redundancy. Extensive experiments show that our approach maintains strong cross-modal understanding while substantially reducing computational overhead and exhibiting broad compatibility, paving the way for efficient and scalable VLLM inference.